Two long, straight wires at the corners of a square carry equal magnitude currents as shown below. What is the direction of the net magnetic field at the square's center (the wires are perpendicular to the plane of the screen and the square is in the plane of the screen)?


1
Expert's answer
2021-04-08T07:33:50-0400

To be given in question

Square two long wire magnitude of current equal 

To be asked in question

Square net magnetic field

And directions

We know that

One wire magnetic field"B=[\frac{\mu_{0}I}{4\pi d}](sin\alpha+sin\beta)"

Where i= current flow wire


d=d/2 (distance is square sides to centre)

"\alpha =45°" =U

"\beta=45°" =V

"B= \frac{\mu_{0}{I}} {\pi d} \frac{1}{√2}"

Due to two wire magnetic field "\ufeff\ufeffB=\\frac {\\mu_{0}I}{\\pi d}\u221a2\ufeff\ufeff"

Four wire of squre magnetic field

"B= \frac{\mu_{0}I} {\pi d} \ 2(√2)"


square center net magneti field direction is

downward (bottom of the page)